OpenStudy (anonymous):

A polynomial p of degree 4 leaves a remainder of 2 when divided by x-1 and a remainder of 42 when divided by x+1. Find the polynomial p given that x-2 is a factor of p, p(-2)=128 and p(0)=16

OpenStudy (freckles):

let's right what we are given... p is degree 4 \[p(x)=ax^4+bx^3+cx^2+d \\ \] p(1)=2 <---p divided by (x-1) leaves remainder 2 p(-1)=42 <---p divided by (x+1) leaves remainder 42 p(2)=0 <---p divided by (x-2) leaves remainder 0 p(-2)=128 p(0)=16

OpenStudy (freckles):

oops missed the x term in my 4th degree polynomial

OpenStudy (freckles):

\[p(x)=ax^4+bx^3+cx^2+dx+e \]*

OpenStudy (freckles):

use p(0)=16 to find e

OpenStudy (freckles):

then use the others to give you a system of equations

OpenStudy (freckles):

I have to be right back

OpenStudy (freckles):

in the meantime show me what you have after using all of the conditions I mentioned above

OpenStudy (anonymous):

I got it.... thank you soooo much.. that was helpful.....

OpenStudy (freckles):

np

OpenStudy (anonymous):

I don't know how to solve it......... plz help

OpenStudy (freckles):

you got e right? e is the easiest to find.. the others one will probably cause more difficulty

OpenStudy (freckles):

p(0)=16 was given replace x with 0

OpenStudy (freckles):

\[p(x)=ax^4+bx^3+cx^2+dx+e \\ p(0)=a(0)^4+b(0)^3+c(0)^2+d(0)+e \\ \text{ but } p(0)=16 \\ 16=0+0+0+0+e \\ 16=e\]

OpenStudy (freckles):

Let me see the other equations you get by pluggin in the conditions given